OpenStudy (anonymous):

Simplify. (x – 3)(x^2 – 4x – 7) A. x^3 – 3x^2 + 12x + 21 B. x^3 – 7x^2 + 19x + 21 C. x^3 – 7x^2 + 5x + 21 D. x^3 + x^2 – 5x – 21

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

this is a type of distributive property :) (a+b)(c+d+e) = ac + ad + ae + bc + bd + be

OpenStudy (anonymous):

I know like all the basic math stuff.. But i have never done the distributive property when there isn't something out of the parentheses

OpenStudy (anonymous):

Oh, Im an idoit.. Lol. I didnt read your full response. xD

jigglypuff314 (jigglypuff314):

lol it's alright :) do you get it now? :3

OpenStudy (anonymous):

Yeah, Im solving it now.

OpenStudy (anonymous):

I have no idea.. I keep getting this really weird answers..

jigglypuff314 (jigglypuff314):

hmmm... (x – 3)(x^2 – 4x – 7) -> (x)(x^2) + (x)(-4x) + (x)(-7) + (-3)(x^2) + (-3)(-4x) + (-3)(-7)

OpenStudy (anonymous):

So i know that the first part is x^3

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

I have no idea..

jigglypuff314 (jigglypuff314):

well, what part are you stuck on? :)

OpenStudy (anonymous):

The second part

OpenStudy (anonymous):

I wanna say it would be -4x.. But i dont know..

jigglypuff314 (jigglypuff314):

well, x * x = x^2 right? so x * -4 * x = -4 * x * x = ?

OpenStudy (anonymous):

-4x^2?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Then the next part is -7x?

jigglypuff314 (jigglypuff314):

yes!

OpenStudy (anonymous):

The next part is -3x^2?

jigglypuff314 (jigglypuff314):

yup ^_^

OpenStudy (anonymous):

The next part is 12x?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

The last part is 21?

jigglypuff314 (jigglypuff314):

yes :) now you can add like terms ^_^

OpenStudy (anonymous):

I just got -2x^7 + 21...

OpenStudy (anonymous):

I dont think thats right..

jigglypuff314 (jigglypuff314):

be careful :) add like term Does Not mean that you can add exponents... for example, x^5 + x^3 can Not equal x^8 but 4x^2 + 3x^2 = 7x^2 Does work :)

OpenStudy (anonymous):

Is it C?

jigglypuff314 (jigglypuff314):

Correct! :)

OpenStudy (anonymous):

Thanks so much!